OpenStudy (anonymous):

Please help!!! graph this function.label the axis of symmetry and the vertex. f(x)=-3x^2+9x-8

hero (hero):

The vertex of the graph is \(\left(-\dfrac{b}{2a}, f\left(-\dfrac{b}{2a}\right)\right)\)

hero (hero):

The axis of symmetry is \(x = -\dfrac{b}{2a}\)